Steering Company has expected overhead costs of $14,400,000. The majority of the overhead costs are incurred providing production support to the direct labor force. Direct labor rates vary from $10 to $20 per hour, and more complex tasks are assigned to more skilled workers who have higher pay rates. Steering projects direct labor costs of $2,880,000 and 360,000 direct labor hours. More complex tasks require proportionally more support than do the less complex tasks.

Each model/size of shoe is produced in a single production run and constitutes a job. During the year, the company produced 8,400 pairs of the model K25, size 8 dress shoe. A total of 10,400 direct labor hours, costing $72,800, were assigned to the job.

(a)

Determine the amount of overhead assigned to the model K25, size 8 dress shoe, using both direct labor hours and direct labor cost as the allocation base
Answer #1

a) Overhead rate = Estimated overhead/Estimated allocation Base

On the basis of Direct labor hour =14400000/360000 = $40 per labor hour

On the basis of direct labor cost = 14400000/2880000 = 500% of direct labor cost

Overhead assigned :

On the basis of labor hours = 10400*40 = $416000

On the basis of direct labor cost = 72800*5 = $364000